# A Survey on Large Language Model based Autonomous Agents

### v5

@@ 2.1.3 Planning Module
$\bullet$ Single-path Reasoning. In this strategy, the final task is decomposed into several intermediate steps. These steps are connected in a cascading manner, with each step leading to only one subsequent step. LLMs follow these steps to achieve the final goal. Specifically, Chain of Thought (CoT) [45] proposes inputting reasoning steps for solving complex problems into the prompt. These steps serve as examples to inspire LLMs to plan and act in a step-by-step manner. In this method, the plans are created based on the inspiration from the examples in the prompts. Zero-shot-CoT [46] enables LLMs to generate task reasoning processes by prompting them with trigger sentences like "think step by step". Unlike CoT, this method does not incorporate reasoning steps as examples in the prompts. Re-Prompting [47] involves checking whether each step meets the necessary prerequisites before generating a plan. If a step fails to meet the prerequisites, it introduces a prerequisite error message and prompts the LLM to regenerate the plan. ReWOO [48] introduces a paradigm of separating plans from external observations, where the agents first generate plans and obtain observations independently, and then combine them together to derive the final results. HuggingGPT [13] first decomposes the task into many sub-goals, and then solves each of them based on Huggingface. Different from CoT and Zero-shot-CoT, which outcome all the reasoning steps in a one-shot manner, ReWOO and HuggingGPT produce the results by accessing LLMs multiply times.
@@ 2.1.4 Action Module
Action Goal: The agent can perform actions with various objectives. Here, we present several representative examples: (1) Task Completion. In this scenario, the agent’s actions are aimed at accomplishing specific tasks, such as crafting an iron pickaxe in Minecraft [38] or completing a function in software development [18]. These actions usually have well-defined objectives, and each action contributes to the completion of the final task. Actions aimed at this type of goal are very common in existing literature. (2) Communication. In this case, the actions are taken to communicate with the other agents or real humans for sharing information or collaboration. For example, the agents in ChatDev [18] may communicate with each other to collectively accomplish software development tasks. In Inner Monologue [61], the agent actively engages in communication with humans and adjusts its action strategies based on human feedback. (3) Environment Exploration. In this example, the agent aims to explore unfamiliar environments to expand its perception and strike a balance between exploring and exploiting. For instance, the agent in Voyager [38] may explore unknown skills in their task completion process, and continually refine the skill execution code based on environment feedback through trial and error.
@@
(1) APIs. Leveraging external APIs to complement and expand action space is a popular paradigm in recent years. For example, HuggingGPT [13] leverages the models on HuggingFace to accomplish complex user tasks.  [66, 67] propose to automatically generate queries to extract relevant content from external web pages when responding to user request. TPTU [67] interfaces with both Python interpreters and LaTeX compilers to execute sophisticated computations such as square roots, factorials and matrix operations. Another type of APIs is the ones that can be directly invoked by LLMs based on natural language or code inputs. For instance, Gorilla [68] is a fine-tuned LLM designed to generate accurate input arguments for API calls and mitigate the issue of hallucination during external API invocations. ToolFormer [15] is an LLM-based tool transformation system that can automatically convert a given tool into another one with different functionalities or formats based on natural language instructions. API-Bank [69] is an LLM-based API recommendation agent that can automatically search and generate appropriate API calls for various programming languages and domains. API-Bank also provides an interactive interface for users to easily modify and execute the generated API calls. ToolBench [14] is an LLM-based tool generation system that can automatically design and implement various practical tools based on natural language requirements. The tools generated by ToolBench include calculators, unit converters, calendars, maps, charts, etc. RestGPT [70] connects LLMs with RESTful APIs, which follow widely accepted standards for web services development, making the resulting program more compatible with real-world applications. TaskMatrix.AI [71] connects LLMs with millions of APIs to support task execution. At its core lies a multimodal conversational foundational model that interacts with users, understands their goals and context, and then produces executable code for particular tasks. All these agents utilize external APIs as their tools, and provide interactive interfaces for users to easily modify and execute the generated or transformed tools.
@@
(3) External Models. Previous studies often utilize external models to expand the range of possible actions. In comparison to APIs, external models typically handle more complex tasks. Each external model may correspond to multiple APIs. For example, to enhance the text retrieval capability, MemoryBank [39] incorporates two language models: one is designed to encode the input text, while the other is responsible for matching the query statements. ViperGPT [74] firstly uses Codex, which is implemented based on language model, to generate Python code from text descriptions, and then executes the code to complete the given tasks. TPTU [67] incorporates various LLMs to accomplish a wide range of language generation tasks such as generating code, producing lyrics, and more. ChemCrow [75] is an LLM-based chemical agent designed to perform tasks in organic synthesis, drug discovery, and material design. It utilizes seventeen expert-designed models to assist its operations. MM-REACT [76] integrates various external models, such as VideoBERT for video summarization, X-decoder for image generation, and SpeechBERT for audio processing, enhancing its capability in diverse multimodal scenarios.
@@
$\bullet$ Internal Knowledge. In addition to utilizing external tools, many agents rely solely on the internal knowledge of LLMs to guide their actions. We now present several crucial capabilities of LLMs that can support the agent to behave reasonably and effectively. (1) Planning Capability. Previous work has demonstrated that LLMs can be used as decent planers to decompose complex task into simpler ones [45]. Such capability of LLMs can be even triggered without incorporating examples in the prompts [46]. Based on the planning capability of LLMs, DEPS [33] develops a Minecraft agent, which can solve complex task via sub-goal decomposition. Similar agents like GITM [16] and Voyager [38] also heavily rely on the planning capability of LLMs to successfully complete different tasks. (2) Conversation Capability. LLMs can usually generate high-quality conversations. This capability enables the agent to behave more like humans. In the previous work, many agents take actions based on the strong conversation capability of LLMs. For example, in ChatDev [18], different agents can discuss the software developing process, and even can make reflections on their own behaviors. In RLP [30], the agent can communicate with the listeners based on their potential feedback on the agent’s utterance. (3) Common Sense Understanding Capability. Another important capability of LLMs is that they can well comprehend human common sense. Based on this capability, many agents can simulate human daily life and make human-like decisions. For example, in Generative Agent, the agent can accurately understand its current state, the surrounding environment, and summarize high-level ideas based on basic observations. Without the common sense understanding capability of LLMs, these behaviors cannot be reliably simulated. Similar conclusions may also apply to RecAgent [21] and S3 [77], where the agents aim to simulate user recommendation and social behaviors.
@@
Action Impact: Action impact refers to the consequences of the action. In fact, the action impact can encompass numerous instances, but for brevity, we only provide a few examples. (1) Changing Environments. Agents can directly alter environment states by actions, such as moving their positions, collecting items, constructing buildings, etc. For instance, in GITM [16] and Voyager [38], the environments are changed by the actions of the agents in their task completion process. For example, if the agent mines three woods, then they may disappear in the environments. (2) Altering Internal States. Actions taken by the agent can also change the agent itself, including updating memories, forming new plans, acquiring novel knowledge, and more. For example, in Generative Agents [20], memory streams are updated after performing actions within the system. SayCan [78] enables agents to take actions to update understandings of the environment. (3) Triggering New Actions. In the task completion process, one agent action can be triggered by another one. For example, Voyager [38] constructs buildings once it has gathered all the necessary resources.
@@ 2.2 Agent Capability Acquisition
In the above sections, we mainly focus on how to design the agent architecture to better inspire the capability of LLMs to make it qualified for accomplishing tasks like humans. The architecture functions as the “hardware” of the agent. However, relying solely on the hardware is insufficient for achieving effective task performance. This is because the agent may lack the necessary task-specific capabilities, skills and experiences, which can be regarded as "software" resources. In order to equip the agent with these resources, various strategies have been devised. Generally, we categorize these strategies into two classes based on whether they require fine-tuning of the LLMs. In the following, we introduce each of them more in detail.
@@
Capability Acquisition with Fine-tuning: A straightforward method to enhance the agent capability for task completion is fine-tuning the agent based on task-dependent datasets. Generally, the datasets can be constructed based on human annotation, LLM generation or collected from real-world applications. In the following, we introduce these methods more in detail.
@@
$\bullet$ Fine-tuning with Human Annotated Datasets. To fine-tune the agent, utilizing human annotated datasets is a versatile approach that can be employed in various application scenarios. In this approach, researchers first design annotation tasks and then recruit workers to complete them. For example, in CoH [84], the authors aim to align LLMs with human values and preferences. Different from the other models, where the human feedback is leveraged in a simple and symbolic manner, this method converts the human feedback into detailed comparison information in the form of natural languages. The LLMs are directly fine-tuned based on these natural language datasets. In RET-LLM [42], in order to better convert natural languages into structured memory information, the authors fine-tune LLMs based on a human constructed dataset, where each sample is a “triplet-natural language” pair. In WebShop [85], the authors collect 1.18 million real-world products form amazon.com, and put them onto a simulated e-commerce website, which contains several carefully designed human shopping scenarios. Based on this website, the authors recruit 13 workers to collect a real-human behavior dataset. At last, three methods based on heuristic rules, imitation learning and reinforcement learning are trained based on this dataset. Although the authors do not fine-tune LLM-based agents, we believe that the dataset proposed in this paper holds immense potential to enhance the capabilities of agents in the field of web shopping. In EduChat [86], the authors aim to enhance the educational functions of LLMs, such as open-domain question answering, essay assessment, Socratic teaching, and emotional support. They fine-tune LLMs based on human annotated datasets that cover various educational scenarios and tasks. These datasets are manually evaluated and curated by psychology experts and frontline teachers. SWIFTSAGE [87] is an agent influenced by the dual-process theory of human cognition [88], which is effective for solving complex interactive reasoning tasks. In this agent, the SWIFT module constitutes a compact encoder-decoder language model, which is fine-tuned using human-annotated datasets.
@@
$\bullet$ Fine-tuning with LLM Generated Datasets. Building human annotated dataset needs to recruit people, which can be costly, especially when one needs to annotate a large amount of samples. Considering that LLMs can achieve human-like capabilities in a wide range of tasks, a natural idea is using LLMs to accomplish the annotation task. While the datasets produced from this method can be not as perfect as the human annotated ones, it is much cheaper, and can be leveraged to generate more samples. For example, in ToolBench [14], to enhance the tool-using capability of open-source LLMs, the authors collect 16,464 real-world APIs spanning 49 categories from the RapidAPI Hub. They used these APIs to prompt ChatGPT to generate diverse instructions, covering both single-tool and multi-tool scenarios. Based on the obtained dataset, the authors fine-tune LLaMA [9], and obtain significant performance improvement in terms of tool using. In [82], to empower the agent with social capability, the authors design a sandbox, and deploy multiple agents to interact with each other. Given a social question, the central agent first generates initial responses. Then, it shares the responses to its nearby agents for collecting their feedback. Based on the feedback as well as its detailed explanations, the central agent revise its initial responses to make them more consistent with social norms. In this process, the authors collect a large amount of agent social interaction data, which is then leveraged to fine-tune the LLMs.
@@
$\bullet$ Fine-tuning with Real-world Datasets. In addition to building datasets based on human or LLM annotation, directly using real-world datasets to fine-tune the agent is also a common strategy. For example, in MIND2WEB [89], the authors collect a large amount of real-world datasets to enhance the agent capability in the web domain. In contrast to prior studies, the dataset presented in this paper encompasses diverse tasks, real-world scenarios, and comprehensive user interaction patterns. Specifically, the authors collect over 2,000 open-ended tasks from 137 real-world websites spanning 31 domains. Using this dataset, the authors fine-tune LLMs to enhance their performance on web-related tasks, including movie discovery and ticket booking, among others. In SQL-PALM [90], researchers fine-tune PaLM-2 based on a cross-domain large-scale text-to-SQL dataset called Spider. The obtained model can achieve significant performance improvement on text-to-SQL tasks.
@@
Capability Acquisition without Fine-tuning: In the era of tradition machine learning, the model capability is mainly acquired by learning from datasets, where the knowledge is encoded into the model parameters. In the era of LLMs, the model capability can be acquired either by training/fine-tuning the model parameters or designing delicate prompts (i.e., prompt engineer). In prompt engineer, one needs to write valuable information into the prompts to enhance the model capability or unleash existing LLM capabilities. In the era of agents, the model capability can be acquired based on three strategies: (1) model fine-tuning, (2) prompt engineer and (3) designing proper agent evolution mechanisms (we called it as mechanism engineering). Mechanism engineering is a broad concept that involves developing specialized modules, introducing novel working rules, and other strategies to enhance agent capabilities. For clearly understanding such transitions on the strategy of model capability acquisition, we illustrate them in Figure 4. In the following, we introduce prompting engineering and mechanism engineering for agent capability acquisition.
$\bullet$ Prompting Engineering. Due to the strong language comprehension capabilities, people can directly interact with LLMs using natural languages. This introduces a novel strategy for enhancing agent capabilities, that is, one can describe the desired capability using natural language and then use it as prompts to influence LLM actions. For example, in CoT [45], in order to empower the agent with the capability for complex task reasoning, the authors present the intermediate reasoning steps as few-shot examples in the prompt. Similar techniques are also used in CoT-SC [49] and ToT [50]. In SocialAGI [30], in order to enhance the agent self-awareness capability in conversation, the authors prompt LLMs with the agent beliefs about the mental states of the listeners and itself, which makes the generated utterance more engaging and adaptive. In addition, the authors also incorporate the target mental states of the listeners, which enables the agents to make more strategic plans. Retroformer [91] presents a retrospective model that enables the agent to generate reflections on its past failures. The reflections are integrated into the prompt of LLMs to guide the agent’s future actions. Additionally, this model utilizes reinforcement learning to iteratively improve the retrospective model, thereby refining the LLM prompt.
@@
(1) Trial-and-error. In this method, the agent first performs an action, and subsequently, a pre-defined critic is invoked to judge the action. If the action is deemed unsatisfactory, then the agent reacts by incorporating the critic’s feedback. In RAH [92], the agent serves as a user assistant in recommender systems. One of the agent’s crucial roles is to simulate human behavior and generate responses on behalf of the user. To fulfill this objective, the agent first generates a predicted response and then compares it with the real human feedback. If the predicted response and the real human feedback differ, the critic generates failure information, which is subsequently incorporated into the agent’s next action. In DEPS [33], the agent first designs a plan to accomplish a given task. In the plan execution process, if an action fails, the explainer generates specific details explaining the cause of the failure. This information is then incorporated by the agent to redesign the plan. In RoCo [93], the agent first proposes a sub-task plan and a path of 3D waypoints for each robot in a multi-robot collaboration task. The plan and waypoints are then validated by a set of environment checks, such as collision detection and inverse kinematics. If any of the checks fail, the feedback is appended to each agent’s prompt and another round of dialog begins. The agents use LLMs to discuss and improve their plan and waypoints until they pass all validations. In PREFER [94], the agent first evaluates its performance on a subset of data. If it fails to solve certain examples, LLMs are leveraged to generate feedback information reflecting on the reasons of the failure. Based on this feedback, the agent improves itself by iteratively refining its actions.
@@
(2) Crowd-sourcing. In [95], the authors design a debating mechanism that leverages the wisdom of crowds to enhance agent capabilities. To begin with, different agents provide separate responses to a given question. If their responses are not consistent, they will be prompted to incorporate the solutions from other agents and provide an updated response. This iterative process continues until reaching a final consensus answer. In this method, the capability of each agent is enhance by understanding and incorporating the other agents’ opinions.
@@
(3) Experience Accumulation. In GITM [16], the agent does not know how to solve a task in the beginning. Then, it makes explorations, and once it has successfully accomplished a task, the actions used in this task are stored into the agent memory. In the future, if the agent encounters a similar task, then the relevant memories are extracted to complete the current task. In this process, the improved agent capability comes from the specially designed memory accumulation and utilization mechanisms. In Voyager [38], the authors equip the agent with a skill library, and each skill in the library is represented by executable codes. In the agent-environment interaction process, the codes for each skill will be iteratively refined according to the environment feedback and the agent self-verification results. After a period of execution, the agent can successfully complete different tasks efficiently by accessing the skill library. In AppAgent [96], the agent is designed to interact with apps in a manner akin to human users, learning through both autonomous exploration and observation of human demonstrations. Throughout this process, it constructs a knowledge base, which serves as a reference for performing intricate tasks across various applications on a mobile phone. In MemPrompt [97], the users are requested to provide feedback in natural language regarding the problem-solving intentions of the agent, and this feedback is stored in memory. When the agent encounters similar tasks, it attempts to retrieve related memories to generate more suitable responses.
@@
(4) Self-driven Evolution. In LMA3 [98], the agent can autonomously set goals for itself, and gradually improve its capability by exploring the environment and receiving feedback from a reward function. Following this mechanism, the agent can acquire knowledge and develop capabilities according to its own preferences. In SALLM-MS [99], by integrating advanced large language models like GPT-4 into a multi-agent system, agents can adapt and perform complex tasks, showcasing advanced communication capabilities, thereby realizing self-driven evolution in their interactions with the environment. In CLMTWA [100], by using a large language model as a teacher and a weaker language model as a student, the teacher can generate and communicate natural language explanations to improve the student’s reasoning skills via theory of mind. The teacher can also personalize its explanations for the student and intervene only when necessary, based on the expected utility of intervention. In NLSOM [101], different agents communicate and collaborate through natural language to solve tasks that a single agent cannot solve. This can be seen as a form of self-driven learning, utilizing the exchange of information and knowledge between multiple agents. However, unlike other models such as LMA3, SALLM-MS, and CLMTWA, NLSOM allows for dynamic adjustment of agent roles, tasks, and relationships based on the task requirements and feedback from other agents or the environment.

### v6

@@ 2.1.3 Planning Module
$\bullet$ Single-path Reasoning. In this strategy, the final task is decomposed into several intermediate steps. These steps are connected in a cascading manner, with each step leading to only one subsequent step. LLMs follow these steps to achieve the final goal. Specifically, Chain of Thought (CoT) [45] proposes inputting reasoning steps for solving complex problems into the prompt. These steps serve as examples to inspire LLMs to plan and act in a step-by-step manner. In this method, the plans are created based on the inspiration from the examples in the prompts. Zero-shot-CoT [46] enables LLMs to generate task reasoning processes by prompting them with trigger sentences like "think step by step". Unlike CoT, this method does not incorporate reasoning steps as examples in the prompts. Re-Prompting [47] involves checking whether each step meets the necessary prerequisites before generating a plan. If a step fails to meet the prerequisites, it introduces a prerequisite error message and prompts the LLM to regenerate the plan. ReWOO [48] introduces a paradigm of separating plans from external observations, where the agents first generate plans and obtain observations independently, and then combine them together to derive the final results. HuggingGPT [13] first decomposes the task into many sub-goals, and then solves each of them based on Huggingface. Different from CoT and Zero-shot-CoT, which outcome all the reasoning steps in a one-shot manner, ReWOO and HuggingGPT produce the results by accessing LLMs multiply times. SWIFTSAGE [49], inspired by the dual-process theory of human cognition [50], combines the power of both SWIFT and SAGE modules for planning in complex interactive tasks. The SWIFT module provides quick responses based on learned patterns, while the SAGE module, using large language models, conducts in-depth planning by asking key questions and generating action sequences to ensure successful task completion.
@@ 2.1.4 Action Module
Action Goal: The agent can perform actions with various objectives. Here, we present several representative examples: (1) Task Completion. In this scenario, the agent’s actions are aimed at accomplishing specific tasks, such as crafting an iron pickaxe in Minecraft [38] or completing a function in software development [18]. These actions usually have well-defined objectives, and each action contributes to the completion of the final task. Actions aimed at this type of goal are very common in existing literature. (2) Communication. In this case, the actions are taken to communicate with the other agents or real humans for sharing information or collaboration. For example, the agents in ChatDev [18] may communicate with each other to collectively accomplish software development tasks. In Inner Monologue [62], the agent actively engages in communication with humans and adjusts its action strategies based on human feedback. (3) Environment Exploration. In this example, the agent aims to explore unfamiliar environments to expand its perception and strike a balance between exploring and exploiting. For instance, the agent in Voyager [38] may explore unknown skills in their task completion process and continually refine the skill execution code based on environment feedback through trial and error.
@@
(1) APIs. Leveraging external APIs to complement and expand action space is a popular paradigm in recent years. For example, HuggingGPT [13] integrates HuggingFace’s vast model ecosystem to tackle complex user tasks. Similarly, WebGPT [67] proposes to automatically generate queries to extract relevant content from external web pages when responding to user request. TPTU [68] explores the potential of LLMs to address intricate tasks through strategic task planning and API-based tools. Gorilla [69] introduces a fine-tuned LLM capable of generating precise input arguments for API calls, effectively mitigating hallucination issues during external API usage. ToolFormer [15] employs self-supervised learning to determine when and how to invoke external tools, using demonstrations of tool APIs for training. API-Bank [70] offers a comprehensive benchmark with a diverse collection of API tools to systematically evaluate tool-augmented LLMs, alongside robust training datasets designed to enhance their integration capabilities. ToolLLaMA [14] proposes a tool-use framework encompassing data collection, training, and evaluation, with the resulting fine-tuned model excelling across a wide array of APIs. RestGPT [71] connects LLMs with RESTful APIs, which follow widely accepted standards for web services development, making the resulting program more compatible with real-world applications. TaskMatrix.AI [72] connects LLMs with an extensive ecosystem of APIs to support task execution. At its core lies a multimodal conversational foundational model that interacts with users, understands their goals and context, and then produces executable code for particular tasks. In essence, these intelligent agents strategically harness external APIs as versatile tools, systematically expanding their action space and transcending the inherent limitations of traditional language models by integrating diverse computational capabilities.
@@
(3) External Models. Previous studies often utilize external models to expand the range of possible actions. In comparison to APIs, external models typically handle more complex tasks. Each external model may correspond to multiple APIs. For example, ViperGPT [75] firstly uses Codex, which is implemented based on language model, to generate Python code from text descriptions, and then executes the code to complete the given tasks. ChemCrow [76] is an LLM-based chemical agent designed to perform tasks in organic synthesis, drug discovery, and material design. It utilizes seventeen expert-designed models to assist its operations. MM-REACT [77] integrates various external models, such as VideoBERT for video summarization, X-decoder for image generation, and SpeechBERT for audio processing, enhancing its capability in diverse multimodal scenarios.
@@
$\bullet$ Internal Knowledge. In addition to utilizing external tools, many agents rely solely on the internal knowledge of LLMs to guide their actions. We now present several crucial capabilities of LLMs that can support the agent to behave reasonably and effectively. (1) Planning Capability. Previous work has demonstrated that LLMs can be used as decent planners to decompose complex tasks into simpler ones [45]. Such a capability of LLMs can be even triggered without incorporating examples in the prompts [46]. Building on the planning capability of LLMs, DEPS [33] develops a Minecraft agent, which can solve complex tasks via sub-goal decomposition. Similar agents like GITM [16] and Voyager [38] also heavily rely on the planning capability of LLMs to successfully complete various tasks. (2) Conversation Capability. LLMs can usually generate high-quality conversations. This capability enables agents to behave more like humans. In the previous work, many agents take actions based on the strong conversation capability of LLMs. For example, in ChatDev [18], different agents can discuss the software development process, and even ceven reflect on their own behaviors. In RLP [30], the agent can communicate with the listeners based on their potential feedback on the agent’s utterance. (3) Common Sense Understanding Capability. Another important capability of LLMs is that they can well comprehend human common sense. Based on this capability, many agents can simulate human daily life and make human-like decisions. For example, in Generative Agent [20], the agent can accurately understand its current state, the surrounding environment, and summarize high-level ideas based on basic observations. Without the common sense understanding capability of LLMs, these behaviors cannot be reliably simulated. Similar conclusions may also apply to RecAgent [21] and S3 [78], where the agents focus on simulating user social behaviors.
@@
Action Impact: Action impact refers to the consequences of an agent’s actions. While the range of possible impacts is vast, we highlight a few key examples for clarity: (1) Changing Environments. Agents can directly alter environment states by actions, such as moving their positions, collecting items, constructing buildings, etc. For instance, in GITM [16] and Voyager [38], the environments are changed by the actions of the agents in their task completion process. Specifically, when an agent collects resources—such as harvesting three pieces of wood—the resources disappear from the environment. (2) Altering Internal States. Actions taken by the agent can also change the agent itself, including updating memories, forming new plans, acquiring novel knowledge, and more. For example, in Generative Agents [20], memory streams are updated after performing actions within the system. Similarly, SayCan [79] enables agents to take actions to update understandings of the environment. (3) Triggering New Actions. In task completion processes, one action often leads to subsequent actions. For example, in Voyager [38], once the agent has gathered the necessary resources, it triggers the construction of buildings.
@@ 2.2 Agent Capability Acquisition
In the sections above, we focus mainly on how to design the agent architecture to better harness the capabilities of LLMs to enabling them to accomplish tasks akin to human performance. The architecture functions as the “hardware” of an agent. However, relying solely on the hardware is insufficient for achieving effective task performance. This is because the agent may lack the necessary task-specific capabilities, skills, and experiences, which can be regarded as "software" resources. In order to equip the agent with these resources, various strategies have been devised. Generally, we categorize these strategies into two classes based on whether they require fine-tuning of the LLMs. Below, we introduce each category in detail.
@@
Capability Acquisition with Fine-tuning: A direct approach to enhance agent capabilities for task completion is to fine-tune the model using task-specific datasets. These datasets can be constructed from human annotations, LLM-generated content, or real-world applications. We discuss these methods in detail below.
@@
$\bullet$ Fine-tuning with Human Annotated Datasets. To fine-tune the agent, utilizing human annotated datasets is a versatile approach that can be employed in various application scenarios. In this approach, researchers first design annotation tasks and then recruit workers to complete them. For example, in CoH [85], the authors aim to align LLMs with human values and preferences. Different from the other models, where the human feedback is leveraged in a simple and symbolic manner, this method converts the human feedback into detailed comparison information in the form of natural languages. The LLMs are directly fine-tuned based on these natural language datasets. In RET-LLM [41], in order to better convert natural languages into structured memory information, the authors fine-tune LLMs based on a human constructed dataset, where each sample is a “triplet-natural language” pair. In WebShop [86], the authors collect 1.18 million real-world products from amazon.com, and put them onto a simulated e-commerce website, which contains several carefully designed human shopping scenarios. Based on this website, the authors recruit 13 workers to collect a real-human behavior dataset. At last, three methods based on heuristic rules, imitation learning and reinforcement learning are trained based on this dataset. Although the authors do not fine-tune LLM-based agents, we believe that the dataset proposed in this paper holds immense potential to enhance the capabilities of agents in the field of web shopping. In EduChat [87], the authors aim to enhance the educational functions of LLMs, such as open-domain question answering, essay assessment, Socratic teaching, and emotional support. They fine-tune LLMs based on human annotated datasets that cover various educational scenarios and tasks.
@@
$\bullet$ Fine-tuning with LLM Generated Datasets. Building human-annotated datasets typically requires recruiting people, which can be costly, especially when dealing with large-scale annotation tasks. Considering that LLMs can achieve human-like capabilities in a wide range of tasks, a natural idea is using LLMs to accomplish the annotation task. While the datasets produced from this method can be not as perfect as the human annotated ones, it is much cheaper, and can be leveraged to generate more samples. For example, in ToolBench [14], to enhance the tool-using capability of open-source LLMs, the authors collect 16,464 real-world APIs spanning 49 categories from the RapidAPI Hub. They used these APIs to prompt ChatGPT to generate diverse instructions, covering both single-tool and multi-tool scenarios. Based on the obtained dataset, the authors fine-tune LLaMA [9], and obtain significant performance improvement in terms of tool using. In [83], to empower the agent with social capability, the authors design a sandbox, and deploy multiple agents to interact with each other. Given a social question, the central agent first generates initial responses. Then, it shares the responses to its nearby agents for collecting their feedback. Based on the feedback as well as its detailed explanations, the central agent revise its initial responses to make them more consistent with social norms. In this process, the authors collect a large amount of agent social interaction data, which is then leveraged to fine-tune the LLMs.
@@
$\bullet$ Fine-tuning with Real-world Datasets. In addition to building datasets based on human or LLM annotations, directly using real-world datasets to fine-tune the agent is also a common strategy. For example, in MIND2WEB [88], the authors collect a large amount of real-world datasets to enhance the agent capability in the web domain. In contrast to prior studies, the dataset presented in this paper encompasses diverse tasks, real-world scenarios, and comprehensive user interaction patterns. Specifically, the authors collect over 2,000 open-ended tasks from 137 real-world websites spanning 31 domains. Using this dataset, the authors fine-tune LLMs to enhance their performance on web-related tasks such as movie discovery and ticket booking. Similarly, in SQL-PaLM [89], researchers fine-tune PaLM-2 using cross-domain, large-scale text-to-SQL datasets, including Spider and BIRD. The resulting model achieves notable performance improvements on text-to-SQL tasks, demonstrating the effectiveness of real-world datasets for domain-specific applications.
@@
Capability Acquisition without Fine-tuning: In the era of tradition machine learning, the model capability is mainly acquired by learning from datasets, where the knowledge is encoded into the model parameters. In the era of LLMs, the model capability can be acquired either by training/fine-tuning the model parameters or designing delicate prompts (i.e., prompt engineering). In prompt engineering, one needs to write valuable information into the prompts to enhance the model capability or unleash existing LLM capabilities. In the era of agents, the model capability can be acquired based on three strategies: (1) model fine-tuning, (2) prompt engineering and (3) designing proper agent evolution mechanisms (we called it as mechanism engineering). Mechanism engineering is a broad concept that involves developing specialized modules, introducing novel working rules, and other strategies to enhance agent capabilities. For clearly understanding the transitions of model capability acquisition strategies, we illustrate them in Figure 4. In the following, we detail prompting engineering and mechanism engineering.
$\bullet$ Prompting Engineering. Due to the strong language comprehension capabilities, people can directly interact with LLMs using natural languages. This introduces a novel strategy for enhancing agent capabilities, that is, one can describe the desired capability using natural language and then use it as prompts to influence LLM actions. For example, in CoT [45], in order to empower the agent with the capability for complex task reasoning, the authors present the intermediate reasoning steps as few-shot examples in the prompt. Similar techniques are also used in CoT-SC [51] and ToT [52]. In RLP [30], the authors aim to enhance an agent’s self-awareness in conversations by prompting LLMs with the agent’s beliefs about both its own and the listeners’ mental states. This approach results in more engaging and adaptive utterances. Furthermore, the incorporation of the target mental states of listeners allows the agent to formulate more strategic plans. Retroformer [90] presents a retrospective model that enables the agent to generate reflections on its past failures. The reflections are integrated into the prompt of LLMs to guide the agent’s future actions. Additionally, this model utilizes reinforcement learning to iteratively improve the retrospective model, thereby refining the LLM prompt.
@@
(1) Trial-and-error. In this method, the agent first performs an action, and subsequently, a pre-defined critic is invoked to judge the action. If the action is deemed unsatisfactory, then the agent reacts by incorporating the critic’s feedback. For example, in RAH [91], the agent serves as a user assistant in recommender systems. One of the agent’s crucial roles is to simulate human behavior and generate responses on behalf of the user. To fulfill this objective, the agent first generates a predicted response and then compares it with the real human feedback. If the predicted response and the real human feedback differ, the critic generates failure information, which is subsequently incorporated into the agent’s next action. Similarly, in DEPS [33], the agent first designs a plan to accomplish a given task. In the plan execution process, if an action fails, the explainer generates specific details explaining the cause of the failure. This information is then incorporated by the agent to redesign the plan. In RoCo [92], the agent first proposes a sub-task plan and a path of 3D waypoints for each robot in a multi-robot collaboration task. The plan and waypoints are then validated by a set of environment checks, such as collision detection and inverse kinematics. If any of the checks fail, the feedback is appended to each agent’s prompt and another round of dialog begins. The agents use LLMs to discuss and improve their plan and waypoints until they pass all validations. PREFER [93] extends this idea by leveraging LLMs to generate detailed feedback when the agent underperforms, enabling iterative refinement and performance improvement.
@@
(2) Crowd-sourcing. In [94], the authors design a debating mechanism that leverages the wisdom of crowds to enhance agent capabilities. To begin with, different agents provide separate responses to a given question. If their responses are not consistent, they will be prompted to incorporate the solutions from other agents and provide an updated response. This iterative process continues until reaching a final consensus answer. In this method, the capability of each agent is enhanced by understanding and incorporating the other agents’ opinions.
@@
(3) Experience Accumulation. In GITM [16], the agent does not know how to solve a task in the beginning. Then, it makes explorations, and once it has successfully accomplished a task, the actions used in this task are stored into the agent memory. In the future, if the agent encounters a similar task, then the relevant memories are extracted to complete the current task. In this process, the improved agent capability comes from the specially designed memory accumulation and utilization mechanisms. Voyager [38] introduces a skill library, where executable codes for specific skills are refined through interactions with the environment, enabling efficient task execution over time. In AppAgent [95], the agent is designed to interact with apps in a manner akin to human users, learning through both autonomous exploration and observation of human demonstrations. Throughout this process, it constructs a knowledge base, which serves as a reference for performing intricate tasks across various applications on a mobile phone. In MemPrompt [96], the users are requested to provide feedback in natural language regarding the problem-solving intentions of the agent, and this feedback is stored in memory. When the agent encounters similar tasks, it attempts to retrieve related memories to generate more suitable responses.
@@
(4) Self-driven Evolution. This method allows agents to autonomously improve through self-directed learning and feedback mechanisms. LMA3 [97] enables the agent to autonomously set goals for itself, and gradually improve its capability by exploring the environment and receiving feedback from a reward function. Following this mechanism, the agent can acquire knowledge and develop capabilities according to its own preferences. SALLM-MS [98] integrates advanced LLMs like GPT-4 into a multi-agent system, agents can adapt and perform complex tasks, showcasing advanced communication capabilities, thereby realizing self-driven evolution in their interactions with the environment. In CLMTWA [99], by using a large language model as a teacher and a weaker language model as a student, the teacher can generate and communicate natural language explanations to improve the student’s reasoning skills via theory of mind. The teacher can also personalize its explanations for the student and intervene only when necessary, based on the expected utility of intervention. Meanwhile, NLSOM [100],leverages natural language collaboration between agents, dynamically adjusting roles, tasks, and relationships based on feedback to solve problems beyond the scope of a single agent.